Find tbe density of air at 0°C and 1.013 x 105 Pa. One mole of air may be assumed to be 79 percent nitrogen (Mnitrogen= 28.0 g) and 21 percent oxygen (Moxygen = 32.0 g).


1
Expert's answer
2021-02-28T07:33:13-0500

"T = 273 \;K \\

P = 1.013 \times 10^5 \;Pa \\

Density = \frac{P \times M}{R\times T} \\

R = 8.314 \times 10^3 \; L\;Pa/K\;mol \\

Density_{nitrogen} = \frac{1.013 \times 10^5 \times 28.0}{8.314 \times 10^3 \times 273} = 1.249 \;g/l \\

Density_{oxygen} = \frac{1.013 \times 10^5 \times 32.0}{8.314 \times 10^3 \times 273} = 1.428 \;g/l \\

Density_{air} = 0.79 \times 1.249 + 0.21 \times 1.428 = 0.9867 + 0.2998 = 1.2865 \;g/l"

Answer: 1.2865 g/l